The product of real roots of the equation 
lxl6/5-26lxl3/5-27=0

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Write,
lxl6/5-26lxl3/5-27=0
as
lxl(3/5)^2 -26lxl3/5-27=0
 
let, |x|^(3/5) = t
the quadratic form will be:
t^2 – 26t + 27 = 0
it has two roots.
t = -1   and  t = 27
 
Now defne t
.i.e.
when x  > 0
t = x^3/5
 
when x
t = – x^3/5
put the value of t serial wise and solve and reject  for the two cases.
U will get two final satisfied roots,
.
Multiply them to get the product.